Find an equation of the line containing the given pair of points. (-5,-2) and (-3,-1). The equation of the line in slope-intercept form is y=
Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Given 2 points example.
Find the equation of the line thru the points (2,1) and (3,5)
---------------------
This is a 2 step process. First find the slope of the line thru the points.
slope, m = diffy/diffx
m = (5-1)/(3-2)
m = 4
---------
Now use y = mx + b with either point to find b, the y-intercept.
y = mx + b
5 = 4*3 + b
b = -7
-------
y = mx + b
y = 4x - 7 is the answer.
------------------------
For further assistance, or to check your work, email me via the thank you note.

Answer by richwmiller(17219)

Just curious.
Have you looked for similar problems before posting?
Have you looked for lessons before posting?